OpenStudy (anonymous):

Write the equation in its equivalent exponential form: log3 81 = x

OpenStudy (anonymous):

\[log_b(y)=x\iff b^x=y\]

OpenStudy (anonymous):

....

OpenStudy (anonymous):

replace \(y\) by \(81\) and \(b\) by \(3\)

OpenStudy (anonymous):

so 3^x=81

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and therefore \(x=4\) since \(3^4=81\)